OpenStudy (anonymous):

using words and symbols describe the sequence. 6,12,18,24

OpenStudy (anonymous):

This sequence represents multiples of 3 by even numbers, starting by 2 up to 8.

OpenStudy (anonymous):

basically its 3(n+2) where n starts at 0

OpenStudy (anonymous):

\[\{3(2^n), \forall (n\in N) \bigwedge (1\le n \le 3) \}\]

OpenStudy (anonymous):

Oh wait, no. That's wrong

OpenStudy (anonymous):

\[\{3*2*n), \forall (n\in N) \bigwedge (1\le n \le 4) \}\]